OpenStudy (anonymous):

PLEASE HELP GEOMETRY. WILL GIVE MEDAL! ABC has the vertices 4(0,0) B(-8.5,3) and C(0,6) , and AXY has the vertices A (0,0), X(-3, -8.5), and Y(-6,0). Which transformation can be used on ABC to show that ABC is congruent to AXY? 1. a translation 3 units to the right and two units up 2. a 90 degree counterclockwise about point A 3. a reflection across the x axis followed by a translation 3 units to the right and 2 units up

OpenStudy (zpupster):

go ahead and graph this-- the answer might be obvious

OpenStudy (anonymous):

i think i finally figured it out it was 90 degrees counterclockwise about point A

OpenStudy (anonymous):

I have more trying to figure out ASA SAS SSS maybe you could help me how to figure out the value of y

OpenStudy (anonymous):

@zpupster can you help me?

OpenStudy (anonymous):

I will answer the question for you.

OpenStudy (anonymous):

If the triangles in the diagram X? 6 11 y

OpenStudy (anonymous):

The two triangles are congruent?

OpenStudy (anonymous):

Yes but how do I fiqure out the value of X?

OpenStudy (anonymous):

Look at the diagram. The triangles are congruent, isn't? So ED=BC. As ED=14, BC=x+3; so 14=x+3. x=11

OpenStudy (anonymous):

omg I can't believe I didn't see this. thank you so much for some reason i find this geo confusing and it shouldn't be. :)

OpenStudy (anonymous):

You are welcome!

OpenStudy (anonymous):

thank u @ zpupster